1. List all the subsets of { 8, 16, 27, 31}
2. Determine the number of subsets of {mom, dad, son, daughter}
3. At MegaSalad, a salad can be ordered with some, all, or none of the following set of ingredients on top of the salad greens: {ham, turkey, chicken, tomato, feta cheese, cheddar cheese, cucumbers, onions, red peppers, hot peppers }. How many different variations are there for ordering a salad?
4. Let U = {q, r, s, t, u, v, w, x, y, z}
A = {q, s, u, w, y}
B = {q, s, y, z}
C = {v, w, x, y, z}
Determine A ∩ (B ∪ C)











Answer by stanbon(75887)

1. List all the subsets of { 8, 16, 27, 31}
---
no element: 1
1 element: 4
2 elements: 6
3 elements: 4
4 elements: 1
Total subsets: 2^4 = 16
------------------------
2. Determine the number of subsets of {mom, dad, son, daughter}
Ans: 2^4 = 16
===========================
3. At MegaSalad, a salad can be ordered with some, all, or none of the following set of ingredients on top of the salad greens: {ham, turkey, chicken, tomato, feta cheese, cheddar cheese, cucumbers, onions, red peppers, hot peppers }. How many different variations are there for ordering a salad?
Ans: 2^10 = 1024
-------------------------
4. Let U = {q, r, s, t, u, v, w, x, y, z}
A = {q, s, u, w, y}
B = {q, s, y, z}
C = {v, w, x, y, z}
Determine A ∩ (B ∪ C)
A intersect (qwyzvx)
= {q,w,y}
======================
Cheers,
Stan H.